undeadknight26 (undeadknight26):

Prove that the solution for the equivalent system is the same as the solution for the original system of equations.

undeadknight26 (undeadknight26):

This is what i have so far: 2. A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 a. Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. 2 x 3 + 7 x 3 = 3 6x + 21y = 3 +-3x -4y +5 3x + 17y = 8 b. Use any method to solve the equivalent system of equations (the new first equation with the original second equation). I used the substitution method and substituted x and y for 3.

undeadknight26 (undeadknight26):

@AkashdeepDeb @shamil98 @jdoe0001

OpenStudy (anonymous):

Did you ever get the answer to part c?

undeadknight26 (undeadknight26):

Yea. one sec :D

undeadknight26 (undeadknight26):

Wait...what test is this?

OpenStudy (anonymous):

I figured it out ! But thanks(: